# CONUS Collaboration Final Conus results on coherent elastic neutrino nucleus scattering at the Brokdorf reactor

### v2

@@ Abstract
indicating a chosen CEvNS signal acceptance of 97% for energies above 0.6 keV. The values of the parameters and can be found in table 3.
| Detector | p1 [eV] | p2 [eV] |
@@
The evolution of the main environmental parameters during reactor on and off is shown in figure 8-left and figure 8-right, for the C1 and C4 detectors respectively. The excluded data set during reactor off is significantly longer for the C1 detector as compared to C2, due to the unstable noise conditions for this detector. The shape of the noise peak arising close to the lower end of the ROI was found to follow a Gaussian distribution and its width (FWHM) was monitored over time. Time periods with variations of the FWHM over 1 eV were removed (red-shadowed area). The stability of the peak width (FWHM) of the peak at 10.37 keV from K-shell electron capture in Ge is also displayed with variations below 2 eV for both detectors.

### v3

@@ Abstract
In figure 7 our background model is compared to the reactor off spectrum in the energy region of interest. The spectrum was corrected for the trigger efficiency. This figures demonstrates the compatibility of our reactor off data with the physical background model. In particular, it shows that there is no significant contribution of microphonic noise events to the spectrum. The differences in the spectral shapes for the individual detectors can be explained by the different contributions of e.g. 210Pb within the cryostat or the additional component after the leakage test to the overall spectrum [19].
In previous publications, we reported a significant contribution of noise events at and around the former analysis threshold of about 300 eVee, partly induced by vibrations of the cryocooler systems [13, 4]. Therefore, the setup was rather sensitive to temperature and cryocooler power variations during Run-1 and Run-2. For the Run-5 dataset discussed in this publication, the noise level was well below the background rate described in our background model and the noise variations can be neglected as compared to the expected signal rates. This major improvement is related to our new DAQ system used in Run-5 and the lower, more stable room temperature. To make sure that higher temperatures in the CONUS room and associate higher values of the cryocooler power do not induce a significant rate increase, we analysed data periods removed from the dataset when the air conditioning system was switched off. In such periods, the relative increase in cryocooler power was an order of magnitude higher as compared to the reactor ON/OFF differences. No significant change in the event rate stability was observed even under these extreme conditions. Therefore, we can exclude that noise variations from fluctuations of the environmental parameters significantly enhance the event rate in the reactor off data, encumbering the observation of a signal excess as speculated in [27].
@@
indicating a chosen CEvNS signal acceptance of 97% for energies above 0.6 keVee. The values of the parameters and can be found in table 3.
| Detector | p1 | p2 |
@@
The evolution of the main environmental parameters during reactor on and off is shown in figure 9-left and figure 9-right, for the C1 and C4 detectors respectively. The excluded data set during reactor off is significantly longer for the C1 detector as compared to C2, due to the unstable noise conditions for this detector. The shape of the noise peak arising close to the lower end of the ROI was found to follow a Gaussian distribution and its width (FWHM) was monitored over time. Time periods with variations of the FWHM over 1 eVee were removed (red-shadowed area). The stability of the peak width (FWHM) of the peak at 10.37 keVee from K-shell electron capture in 71Ge is also displayed with variations below 2 eVee for both detectors.
@@
Summary
In this appendix, we provided details about the background model of the CONUS Run-5 dataset. It was obtained using a new DAQ system with highly reduced electronic noise contributions and more sophisticated trigger algorithms as in previous datasets. We could demonstrate that our physical background model explains our reactor off data and that there is no room for an additional significant component from noise events or unknown origin. We explicitly described the fit functions of the quenching models used in this Run-5 analysis and made them available to the community. The results of the individual likelihood fits for the single detectors and the associated spectra are fully consistent. An emphasis during the Run-5 data taking was control and stability of the environmental parameters, which allowed to reduce microphonic noise originating mainly from the detector cryocoolers.